use the following information.
On one of the banks of the Chicago River, there is a water cannon, called the
Water Arc, that sprays recirculated water across the river. The path of the Water
Arc is given by the model
y = º0.006x2 + 1.2x + 10
where x is the distance (in feet) across the river, y is the height of the arc (in feet),
and 10 is the number of feet the cannon is above the river.
What is the maximum height of the water sprayed from the Water Arc?
HELP PLEASE! i'm very stuck on this, and i need to learn it for my test.

Answer by stanbon(75887)

use the following information.
On one of the banks of the Chicago River, there is a water cannon, called the
Water Arc, that sprays recirculated water across the river. The path of the Water
Arc is given by the model
y = -0.006x^2 + 1.2x + 10
where x is the distance (in feet) across the river, y is the height of the arc (in feet),
and 10 is the number of feet the cannon is above the river.
What is the maximum height of the water sprayed from the Water Arc?
-----------------
max occurs when x = -b/(2a) = -1.2/(2*-0.006) = 100
f(100) = -0.006(100)^2 + 1.2*100 + 10 = 70 feet
---
Maximum height: 70 feet
==========================
Cheers,
Stan H.
================